OpenStudy (anonymous):

whats the quartic function with these 2 real zeros given x=3 and x=5

OpenStudy (anonymous):

the zeros are x=3 and x=5 so the factors are, x-3 and x-5 so the function is=(x-3)(x-5)=x^2-3x-5x+15=x^2-8x+15

OpenStudy (anonymous):

i got something different was there a mistake on my end? my answer: y= x^4+ 8x^3 - 16x^2 +8x - 15

OpenStudy (anonymous):

oh sorry i thought quadratic. but it was quartic. so the new right function will be, [(x-3)(x-5)]^2 =[x^2-8x+15]^2 =x^4-16x^3+64x^2+30x^2-240x+225 =x^4-16x^3+94x^2-240x+225